If 25.1mL of 0.152M K3AsO4 (aq) is mixed with 35.3mL of 0.0786M AgNO3 (aq), how many grams of silver arsenate will be produced?
1
Expert's answer
2018-03-09T10:02:07-0500
3AgNO3+ K3AsO4= Ag3AsO4+ KNO3
n(AgNO3)=C*V=0.0786*0.0353=2.778*10^-3 mol
n(K3AsO4)=C*V=0.152*0.0251=3.82*10^-3 mol
n(AgNO3)< n(K3AsO4), AgNO3-limiting reagent
n(Ag3AsO4)= n(AgNO3)/3=2.778*10^-3/3=9.24*10^-4 mol
n(Ag3AsO4)=n*M=9.24*10^-4*419=0.388 g